What is the solution to x/3+(2x+4)/4=(x-1)/6-(3-2x)/2
Answer by praseena(37)

Here we have to solve the following problem,
x/3+(2x+4)/4= (x-1)/6 -(3-2x)
Multiply both sides by the lowest common multiple of 3,4,6,2, which is 12
12(x/3)+12(2x+4)/4=12(x-1)/6 -12(3-2x)/2
now simplify.
Then it comes like this
4x + 3(2x+4) = 2(x-1)-6(3-2x)
4x + 6x + 12 = 2x -2 - 18+ 12x
10x + 12 = 14x - 20
subtract 10x from both sides
then it is 12 = 4x - 20
add 20 to both sides
32 = 4x
divide both sides by 4
then it comes, x = 32/4 = 8
so the solution is x = 8